OpenStudy (vanz):

Solve this equation by factoring or completing the square x^2-10x+21=0

OpenStudy (anonymous):

(x-7)(x-3)

hero (hero):

x^2 - 10x + 21 = 0 x^2-7x-3x + 21 = 0 x(x-7)-3(x-7) = 0 (x-7)(x-3) = 0 x - 7 = 0 x = 7 x -3 = 0 x = 3 x = 3,7

OpenStudy (anonymous):

yeah. . . you identify the roots..